Find the domain and range and sketch the graph of the function f(x)= sqrt 16+6x-x2

1
Expert's answer
2021-11-01T19:36:06-0400

Let us find the domain and range of the function "f(x)= \sqrt {16+6x-x^2}."

Taking into account that

"f(x)= \sqrt {16+6x-x^2}=f(x)= \sqrt {25-(9-6x+x^2)}= \sqrt {25-(3-x)^2},"

we conclude domain contains all points "x" such that "25-(3-x)^2\ge0," which is equivalent to "(3-x)^2\le 25." It follows that "|x-3|\le 5," and hence "-5\le x-3\le 5." We conclude that "-2\le x\le 8," and consequently, the domain of "f" is "[-2,8]." It follows from definition of "f(x)= \sqrt {25-(3-x)^2}" that the maximum of "f" is at the point "x=3," that is the maximum is equal to "5." The minimum is at the point "x=8." Therefore, the range of "f" is "[0,5]."


Let us sketch the graph of the function "f(x)= \sqrt {16+6x-x^2}:"